what is the wavelength of blue light that has a frequency of 6.45 x 10^14 s^-1? (c = 3.00 x 10^8 m/s)
1
Expert's answer
2020-04-02T08:56:21-0400

Wavelength "\lambda" is speed c decided by frequency "\nu" :

"\lambda=c/\nu" = 3*10^8/(6.45*10^14) = 465*10^-9 m = 465 nm